OpenStudy (anonymous):

Solve. –3(5 – 15n) = 15(3n – 2) A. 30 B. –15 C. no solution D. all real numbers

OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2) Expand the left and right side: (-3)(5) + (-3)(-15n) = (15)(3n) + (15)(-2) Tag me if more help needed

OpenStudy (anonymous):

it is no solution because 3(-15n + 5) = 15(3n - 2) then -15(3n - 1) = 15(3n - 2) so you can see that it is no solution because you cannot have 0 = something.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw